the solution need to be in terms of sin and cos(1 point) Consider the linear system -3 -2 = r. 5 3 Find the eigenvalues and eigenvectors for the coefficient matrix. 1 11 =
Answer #1

x IC A 5 3 x=-30-20, x(0)=2 a = 5+ 3X (0) = -5 > for eigen values, (A-111=0 and o de ce, do 5 3-d for eigen vectors, A-df)x=0multiplying equi ② with ~(3ti) and then adding ③ we get, G-i) 6 (3+i) G = 10-2(372) - - 2i6= 2 [5-3 -2] G = è (Qri) = 1 + 21